# CJST: CTC Compressor based Joint Speech and Text Training for Decoder-Only ASR

### v1

@@ I Introduction & Related Work
One common way of feeding continuous acoustic embeddings into a decoder-only model is via an adaptor. A simple and widely-used adaptor is a time reduction layer followed by a linear projection layer [5, 10]. Another approach is to use connectionist temporal classification (CTC) [11] to compress the sequence length based on CTC probabilities, which is referred to as CTC compressor. The concept of CTC compressor was originally proposed for speech translation [12] using attention-based encoder-decoder models [13, 14]. An additional CTC layer was introduced in the encoder to compress the sequence length by averaging neighbouring frames of the same CTC predictions. This idea was further explored in [6] for speech translation using decoder-only models, where the compression can also be done by removing blank frames. Various extensions of the CTC compressor were also explored, e.g., for fully formatted ASR transcription [15] and for streaming ASR [7].
@@
In this work, we propose a novel CTC compressor based joint speech and text training (CJST) framework for decoder-only ASR. More specifically, we propose to utilize the CTC compressor to match speech and text representations from both directions. This is done by combining a simple modality adaptor with several features of the CTC compressor, including sequence compression, on-the-fly forced peaky alignment and CTC class embeddings. Without the need of duration consideration, the proposed CJST achieves effective text injection for both in-domain and corss-domain scenarios, leading to consistent improvements. We also provide a comprehensive study of the CTC compressor, covering various compression modes, edge case handling and behavior under both clean and noisy data conditions, which reveals the most robust setting to use CTC compressor for decoder-only models.
@@ II-A Compression Modes
    Same Prediction Average: Based on the greedy CTC predictions, neighbouring frames of the same predictions are averaged [12, 6].
@@ III-A LM-like
As decoder-only models are naturally language models (LM), the most straightforward way to perform text injection is to train the model as a LM, i.e., no audio prompt at all. In the context of ASR, this can be viewed as internal LM (ILM) training [23, 24] for decoder-only models. We use this LM-like training as baseline for comparison.
@@ IV-B Experiment Settings
For both datasets, we pretrain the audio encoder with CTC loss for 200k steps. We then train the full model (Fig. 1) with corss-entropy (CE) loss, where an auxiliary CTC loss of weight 0.5 is applied to the CTC compressor. The CE training is performed for 200k steps for Librispeech and 500k steps for the in-house data. SpecAugment [32] is always applied. For joint speech and text training, both speech- and text-based CE losses have a weight 1.0. For CJST, we use a simple conformer layer for the modality adaptor, which contains single attention head, convolution kernel 3 and no dimension uplifting for the feed-forward module. It is from-scratch trained with the MSE loss of weight 1.0. We initialize $R_{\text{len}}(h^{\prime},y)$ with 1.0.
@@ IV-D Joint Speech and Text Training
The second experiment set aims to simulate real-world scenarios. Given a seed model trained with sufficient paired speech-text data, we continue training it by injecting both in-domain and out-of-domain text-only data. The goal is to maintain the same (or even better) in-domain performance while achieving a performance boost on the cross-domain evaluation. In this case, we keep 20% paired speech-text data, 30% in-domain text-only data and 50% out-of-domain text-only data for the training. This is performed with the Librispeech audio and LM training data as well as the TED-LIUM2 LM text data, correspondingly. We continue training the Librispeech base models from Table I for another 60k steps. To verify the improvements are not from longer training, we also include results of continuing training the basline with paired speech-text data. The WER results on both Librispeech and TED-LIUM2 test sets are shown in Table IV. Again, we see that the LM-like text injection is already quite effective, while our proposed CJST consistently achieve the best performance in all cases. This matches the goal with about 6% relative improvement on the cross-domain evaluation and small further improvement on the in-domain evaluation.
@@ V Conclusion
In this work, we presented CJST, a novel CTC compressor-based joint speech and text training framework for decoder-only ASR. With a comprehensive evaluation, we firstly showed that CTC compressor can be sensitive under certain compression modes, while the blank probability removal with a high threshold robustly gives the best performance for both clean and noisy data. We then showed that our proposed CJST enables effective joint speech and text training by matching speech and text modalities from both directions. Experiments on both Librispeech and TED-LIUM2 corpora show that CJST achieves the best performance for both in-domain and cross-domain text injection.

### v2

@@ I Introduction & Related Work
One common way of feeding continuous acoustic embeddings into a decoder-only model is via an adaptor. A simple and widely-used adaptor is a time reduction layer followed by a linear projection layer [5, 10]. Another approach is to use connectionist temporal classification (CTC) [11] to compress the sequence length based on CTC probabilities, which is referred to as CTC compressor. The concept of CTC compressor was originally proposed for speech translation [12] using attention-based encoder-decoder models [13, 14]. An additional CTC layer was introduced in the encoder to compress the sequence length by averaging neighboring frames of the same CTC predictions. This idea was further explored in [6] for speech translation using decoder-only models, where the compression can also be done by removing blank frames. Various extensions of the CTC compressor were also explored, e.g., for fully formatted ASR transcription [15] and for streaming ASR [7].
@@
In this work, we propose a novel CTC compressor based joint speech and text training (CJST) framework for decoder-only ASR. More specifically, we propose to utilize the CTC compressor to match speech and text representations from both directions. This is done by combining a simple modality adaptor with several features of the CTC compressor, including sequence compression, on-the-fly forced peaky alignment and CTC class embeddings. Without the need of duration consideration, the proposed CJST achieves effective text injection for both in-domain and cross-domain scenarios, leading to consistent improvements. We also provide a comprehensive study of the CTC compressor, covering various compression modes, edge case handling and behavior under both clean and noisy data conditions, which reveals the most robust setting to use CTC compressor for decoder-only models.
@@ II-A Compression Modes
    Same Prediction Average: Based on the greedy CTC predictions, neighboring frames of the same predictions are averaged [12, 6].
@@ III-A LM-like
As decoder-only models are naturally language models (LM), the most straightforward way to perform text injection is to train the model on text-only data as a LM, i.e., no audio prompt at all. In the context of ASR, this can be viewed as internal LM (ILM) training [23, 24] for decoder-only models. We show that this LM-like training is already quite effective, which is used as a strong baseline for comparison.
@@ IV-B Experiment Settings
For both datasets, we pretrain the audio encoder with CTC loss for 200k steps. We then train the full model (Fig. 1) with cross-entropy (CE) loss, where an auxiliary CTC loss of weight 0.5 is applied to the CTC compressor. The CE training is performed for 200k steps for Librispeech and 500k steps for the in-house data. SpecAugment [32] is always applied. For joint speech and text training, both speech- and text-based CE losses have a weight 1.0. For CJST, we use a simple conformer layer for the modality adaptor, which contains single attention head, convolution kernel 3 and no dimension uplifting for the feed-forward module. It is from-scratch trained with the MSE loss of weight 1.0. We initialize $R_{\text{len}}(h^{\prime},y)$ with 1.0.
@@ IV-D Joint Speech and Text Training
The second experiment set aims to simulate real-world scenarios. Given a seed model trained with sufficient paired speech-text data, we continue training it by injecting both in-domain and out-of-domain text-only data. The goal is to maintain the same (or even better) in-domain performance while achieving a performance boost on the cross-domain evaluation. In this case, we keep 20% paired speech-text data, 30% in-domain text-only data and 50% out-of-domain text-only data for the training. This is performed with the Librispeech audio and LM training data as well as the TED-LIUM2 LM text data, correspondingly. We continue training the Librispeech base models from Table I for another 60k steps. To verify the improvements are not from longer training, we also include results of continuing training the baseline with paired speech-text data. The WER results on both Librispeech and TED-LIUM2 test sets are shown in Table IV. Again, we see that the LM-like text injection is already quite effective, while our proposed CJST consistently achieves further improvements in all cases, fulfilling the goal of this experiment.
Comparing to the baseline trained on paired data only, CJST achieves about 6% relative improvements for both in-domain (Table III) and cross-domain (Table IV) scenarios.
@@ V Conclusion
In this work, we presented CJST, a novel CTC compressor-based joint speech and text training framework for decoder-only ASR. With a comprehensive evaluation, we firstly showed that CTC compressor can be sensitive under certain compression modes, while the blank probability removal with a high threshold robustly gives the best performance for both clean and noisy data. We then showed that LM-like text injection is already quite effective for decoder-only models, while our proposed CJST consistently achieves further improvements by matching speech and text modalities from both directions. Experiments on both Librispeech and TED-LIUM2 corpora show that CJST achieves the best performance for both in-domain and cross-domain text injection.